The demand for a monthly cell phone plans is given by QD = 50 – 0.5 P and the supply of cell phone plans is given by QS = - 25 + P. The quantity is measured in thousands of cell phone plans and the price is in dollars. a. Solve for the equilibrium price and equilibrium quantity of cell phone plans sold in this market. (4) b. Assume that the market price is $ 45 per plan. Would you have a surplus or shortage at that price? Calculate the surplus or shortage.
Answer #1

Af equilibrium a) Bo=0 50-0.5 P = -25+ P 1. 5 p = T 75 P = $50 Equilibrium priu 2 -25+ P - - 25+ 50 - las units - Ea. Quantit
Answer #2
Add you answer here...